OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation x - 4y = 4. slope = 4; y intercept at (0, -1) slope = ; y intercept at (0, 1) slope = ; y intercept at (0, -1) slope = -4; y intercept at (0, 1)

OpenStudy (goformit100):

slope = ; y intercept at (0, -1)

OpenStudy (anonymous):

Rewrite the equation in the form: y = mx + c So, x - 4y = 4 firstly subtract x both the sides..

OpenStudy (anonymous):

Can you do that??

OpenStudy (anonymous):

No problem: so: -4y = -x + 4 multiply -1: 4y = x -4 divide by 4: y = \(\large \frac{1}{4}x - 1\) So the slope is x coefficient that is \(\large \frac{1}{4}\) and y intercept is : (0, -1)

OpenStudy (anonymous):

Thankyou so much